OpenStudy (anonymous):

can anyone explain how to do addiction of 2x and 7x

OpenStudy (anonymous):

i dont get wickerpedia

OpenStudy (anonymous):

like when i entered addiction it said some word like alcol

OpenStudy (anonymous):

So basically: 2x + 7x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Add them together... 2x + 7x = 9x Since they are like terms you can add them like usual... don't forget to keep the variable.

OpenStudy (anonymous):

plz tell me how to do it

OpenStudy (anonymous):

ur wrong the anser given is 14x2

OpenStudy (anonymous):

so its 14x2 so 28

OpenStudy (unklerhaukus):

\[2x+7x=(2+7)x=9x\]

OpenStudy (anonymous):

your wrong

OpenStudy (unklerhaukus):

i was just messing with ya

OpenStudy (anonymous):

okay next time learn matemathicks

OpenStudy (anonymous):

Learn to spell. Then we will learn mathematics.

OpenStudy (anonymous):

i one speling bee last ear

OpenStudy (anonymous):

Of course you did.

OpenStudy (unklerhaukus):

umprestive